OpenStudy (anonymous):

Evaluate -4(a+b)-10a/b when a=8 and b=5 can someone help me with the break down of the problem not the answer just want to know how to work the problem.

hero (hero):

Directly substitute the values of a and b into the expression, then evaluate the expression.

OpenStudy (anonymous):

but how do I evaluate the problem after not sure what I'm evaluating

OpenStudy (anonymous):

alright well u distribute the -4 to the parenthesis, and you put the numbers they give you in the place of a and b

hero (hero):

I assume you have: \[-4(a + b) - 10\frac{a}{b}\] Or is it this: \[-4(a + b) - \frac{10a}{b}\]

hero (hero):

They're both the same anyway, so in either case, plugging in the given values for \(a\) and \(b\) produces: \[-4(8 + 5) - 10\left(\frac{8}{5}\right)\] All you have to do now is evaluate the expression.

OpenStudy (anonymous):

-4( a+b)-10 a/b when a=8 and b=5 that's how it's made out on my worksheet

hero (hero):

So of the expressions I posted above, which one represents what's on your worksheet? To be honest, it doesn't matter because they are both the same.

OpenStudy (anonymous):

\[-4(8+5)-10\left( \frac{ 8 }{ 5 } \right)\]

OpenStudy (anonymous):

thanks @Hero

hero (hero):

Did you computer the correct answer? If so, what is it?

OpenStudy (anonymous):

-68 am I right?

hero (hero):

-4(a + b) -10(a/b) = -4(8 + 5) - 10(8/5) = -(4(13) + 80/5) = -(52 + 16) = -(68) = -68

OpenStudy (anonymous):

thank u :)

hero (hero):

yw